# AttentionRAG: Attention-Guided Context Pruning in Retrieval-Augmented Generation

### v1

@@ 7 Conclusion
In this paper, we propose AttentionRAG, a novel attention-guided context pruning method for RAG systems. The core part of our method is the formatted attention focus mechanism, which constructs an answer hint prefix in a next-token-prediction format for each query, guiding the LLM to pay attention to relevant tokens in the retrieved context through one token. We conduct extensive experiments on the LongBench and Babilong benchmarks, demonstrating its strong performance.
@@
Although we introduce the answer hint prefix paradigm, which can be effectively applied to a wide range of queries, there remain complex queries that it cannot fully address. For example, queries containing multiple sub-questions cannot be easily handled by locating a single focal token through the hint prefix. We propose that future work could explore query decomposition into smaller, unit queries, allowing for separate compression of each part.
@@ Appendix A Pseudocode
15:         Obtain Attention Features $A_{1}$ with $a_{1}$ and $c_{j}$triangleright Eq.  (2)
16:         Get compressed $c_{j}^{\prime}$ according to $A_{1}$ and $c_{j}$triangleright Eq.  (3)
@@ B.1 Prompt for generating answer prefix hint
B.2 Prompt for generating anchor token

### v2

@@ 7 Conclusion
In this paper, we propose AttentionRAG, a novel attention-guided context pruning method for RAG systems. The core part of our method is the formatted attention focus mechanism, which constructs an answer hint prefix and utilize a fixed hint prefix in a next-token-prediction format for each query, guiding the LLM to attend relevant tokens in the retrieved context through one token. We conduct extensive experiments on the 2WikiMQA, HotpotQA, TriviaQA, GovReport and Babilong benchmarks, demonstrating its strong performance.
@@ Appendix A Pseudocode
15:   Obtain Attention Features $A_{1}$ with $a_{1}$ and $c_{j}$ $\triangleright$ Eq.  (2)
16:   Get compressed $c_{j}^{\prime}$ according to $A_{1}$ and $c_{j}$ $\triangleright$ Eq.  (3)
@@ B.1 Prompt for generating answer prefix hint
B.2 Fixed Hint Prefix
Our fixed hint prefix is: "Please output the most relevant keyword or phrase that is relevant to the answer of the question."
B.3 Prompt for generating anchor token
@@ B.4 Hyperparameter settings
B.5 Attention Layer Choice
In LLMs, attention layers capture different levels of information—shallow layers focus on syntax, while deeper ones encode semantics (Ben-Artzy and Schwartz, 2024; Jin et al., 2025). To fully exploit this hierarchy, we aggregate attention scores across all layers for compression, as detailed in §3.2. This mitigates layer-specific bias and captures a broader information spectrum. We compare this approach with using shallow, middle, or deep layers alone on HotpotQA with Llama-3.1-8B-Instruct.
As shown in Table 9, full-layer aggregation yields superior performance, validating our strategy.
[…]
0 - 10 0.35 0.43 4.5x
11 - 20 0.38 0.50 3.6x
21 - 31 0.40 0.48 3.7x
0 - 31 0.42 0.54 3.6x
Table 9: Performance on HotpotQA with different subset of layers